OpenStudy (anonymous):

Which expression is equivalent to sec(–θ)? sec θ -sec θ 1/-cosθ csc θ

OpenStudy (p0sitr0n):

sec(x)=1/cos(x) you know that cos(-x)=cos(x) (cos-pi=cospi, cospi/2=cos-pi/2) so sec(-x)=1/cos(x)=sec(x)

OpenStudy (anonymous):

so whats the answer?

OpenStudy (p0sitr0n):

if you were following the proof, it would be the last line, thus sec(x)

OpenStudy (p0sitr0n):

technically im not even supposed to give the answer right away, just help you get to it